OpenStudy (anonymous):

If you help, I'll give you a medal :D Verify the identity. cot(x- pi/2) = -tan x

OpenStudy (anonymous):

thats easy i will help

OpenStudy (anonymous):

Thank you sooooooo much!!

OpenStudy (anonymous):

cot(x-y)=cotx.coty+1/coty-cotx

OpenStudy (anonymous):

put x=x and y=pi/2

OpenStudy (anonymous):

this gives \[\cot(x-\pi/2)=cotx.\cot(\pi/2)/\cot(\pi/2)-cotx\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

there is a '+1' in the numerator i forgot to type

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

cot pi/2=0

OpenStudy (anonymous):

this gives 1/0-cotx=-tanx

OpenStudy (anonymous):

hence the proof

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

need more explanation??

OpenStudy (anonymous):

I think I got it! Thank you!!!

OpenStudy (anonymous):

be sure..

OpenStudy (anonymous):

lol any doubts...ask me:)

OpenStudy (anonymous):

ok, thank you :D

OpenStudy (anonymous):

my pleasure:D